# The impact of non-Gaussianity when searching for Primordial Black Holes with LISA

### v2

@@ 3.2 The LISA case
$h^{2}\Omega_{\rm GW}(f,A)=A^{2}\Omega_{\rm GW,0}^{\rm Gauss}(f)+A^{3}f_{\rm NL}^{2}h^{2}\Omega^{\rm NG}_{\rm GW,0}(f)$ (3.3)
@@
Finally, consider the amplitudes $A$ corresponding to $f_{\rm PBH}=1$ shown by the dot-dashed lines in Fig. 4. All scenarios above these lines are excluded by PBH overproduction. Crucially, we can observe how, unlike SIGW, NG significantly impacts the abundance of PBH. For example, the required amplitude $A$ for PBH overproduction is reduced by almost an order of magnitude when moving from the Gaussian case $f_{\rm NL}=0$ to $f_{\rm NL}=50$. This effect is especially notable for the wider spectra $\Delta=0.9$, in which case LISA’s ability to prove viable PBH scenarios ($f_{\rm PBH}<1$) is drastically reduced from $f_{\rm NL}=0$ to $f_{\rm NL}=50$. Moreover, as seen in Fig. 2, the suppression of PBH production due to negative NGs is strongest when $f_{\rm NL}\approx-2$, and thus the required amplitude to reach $f_{\rm PBH}=1$ begins to decrease again when $f_{\rm NL}\lesssim-2$. We will return to the implications for PBHs in Sec. 4.

### v3

@@ 3.2 The LISA case
$h^{2}\Omega_{\rm GW}(f,A)=A^{2}h^{2}\Omega_{\rm GW,0}^{\rm Gauss}(f)+A^{3}f_{\rm NL}^{2}h^{2}\Omega^{\rm NG}_{\rm GW,0}(f)$ (3.3)
@@
Finally, consider the amplitudes $A$ corresponding to $f_{\rm PBH}=1$ shown by the dot-dashed lines in Fig. 4. All scenarios above these lines are excluded by PBH overproduction. Crucially, we can observe how, unlike SIGW, NG significantly impacts the abundance of PBH. For example, the required amplitude $A$ for PBH overproduction is reduced by almost an order of magnitude when moving from the Gaussian case $f_{\rm NL}=0$ to $f_{\rm NL}=50$. This effect is especially notable for the wider spectra $\Delta=0.9$, in which case LISA’s ability to prove viable PBH scenarios ($f_{\rm PBH}<1$) is drastically reduced from $f_{\rm NL}=0$ to $f_{\rm NL}=50$. Moreover, as seen in Fig. 2, the suppression of PBH production due to negative NGs is strongest when $f_{\rm NL}\approx-2$, and thus the required amplitude to reach $f_{\rm PBH}=1$ begins to decrease again when $f_{\rm NL}\lesssim-2$. We added the corresponding line in the Figure as a reference. We will return to the implications for PBHs in Sec. 4.